### v1

Abstract
In the context of interstellar communication, it is not known beforehand if the receiver of a given signal would be a plant, an insect, or even life forms unknown to terrestrial scientists. Regardless of the situation, the message time scale could be too fast or too slow and those beings would probably never decode it. Therefore, it is of interest to devise a way to encode messages agnostic of time scale. Fractal messaging would allow one to do this due to their structural self-similarity and, sometimes, scale invariance. By starting from a spatial embedding rationale, a framework is developed for a time scale-free messaging alternative. When one considers a time-agnostic framework for message transmission, it would be interesting to encode a message such that it could be decoded along several spatio-temporal scales. This way, the core idea of the framework hereby proposed is to encode a binary message as waves along infinitely many (power-like distributed) frequencies and amplitudes, transmit such message, and then decode and reproduce it. To do so, the components of the Weierstrass function, a known fractal, are used as the carriers of the message. Each component will have its amplitude modulated to embed the binary stream, allowing for a time-agnostic approach to messaging.
keywords:
Fractal messaging, scale-free communication, space-time agnostic, Weierstrass function, amplitude modulation
1 Introduction
Despite the multitude of formal definitions, the word ‘communicate’ is historically related to the word ‘common’. From an etymological point of view, such historical understanding is validated by noticing that the origins of the word ‘communicate’ trace back to the Latin verb communicare, which means ‘to share’, ‘to make common’, and which is related to the Latin word communis, that translates as common. Therefore, there is a somewhat general agreement that when someone or something communicates, one makes things common [1]. There are no pre-established actors in the process, such that communication may take place between entities of very different complexities.
In order to further explore this aspect of making things common, it is important to study the communication process scientifically, such that its elements and acting entities become characterized. Thus, in order to provide a theoretical structure for the analysis of the communication process, the so-called models of communication are used as conceptual representations which aim to provide a simplified overview of the main components involved in the process of communication [2]. Normally, it is assumed that two entities–the sender and the receiver, exchange something, a signal–that could be meaningful to both. It is true, on the other hand, that the sender and the receiver do not need to be an individual or single entity, they may represent groups. Nevertheless, it is often understood that both the sender and the receiver are known.
In the context of making things common, it is crucial for the communication process that both the sender and the receiver are able to give meaning to the signal. Besides, the signal itself must be perceivable to both these entities. Regarding the first topic, whenever the sender and receiver are known, some general correspondence between the signal and its meaning can be granted. On the other hand, when only the sender is known, in order to try to establish any kind of comprehension, the means used to transmit the signal and the signal itself should be the simplest possible, using basic elements such as electromagnetic waves, as noticed by pioneering works on interstellar communications [3].
Transmitting binary signals has been used to send messages to space in the expectation that they could be decoded and understood by receivers in other parts of the universe. One such example is the Arecibo message sent to space in the 1970s. Recently, an updated, binary-coded message has been developed [4]. These messages may include basic mathematical and physical concepts, enabling the establishment of a universal means of communication. They may contain information on the biochemical composition of life on Earth, the Solar System’s time-stamped position in the Milky Way relative to known globular clusters, as well as digitised depictions of the Solar System, and Earth’s surface. More recently, we published work on a method based on the principles of (algorithmic) information theory to agnostically decode properties of these messages in an optimal and universal fashion allowing a naive receiver to infer the geometric properties of the message, a first step in any attempt to understanding [5]. In another piece of work, we also connected the presence of life and possibly intelligence to the same principles of (algorithmic) information with the former as empirical evidence of low randomness and high structure apparent in the universe [6].
While our previous work covered a specific challenge on the receiving end, the sender has to make multiple assumptions and also faces some challenges. For a signal to be perceived by an entity even before attempting any decoding or interpretation, the signal should be within its dimensional spatio-temporal scales. There has been considerable work on studying the exchange of information between various organisms, most of which can be compiled under the field of biocommunication [7]. During interspecies communications, it may be challenging to precisely characterize the means used to transmit a given message, since different species have particular biostructures which enable (or disable) the adequate flow of signals. For example, plants have rigid cell walls which restrict movement and usually prevent plants from sending and receiving signals that depend on rapid movement [8, 9].
Movement is deeply related to space and time, such that communication may be prevented if both sender and receiver function in different spatio-temporal scales. Let one imagine a message is sent to the space. It is not known beforehand if the receiver would be a plant, an insect or even life forms totally unknown to terrestrial scientists. Regardless of the situation, the message time scale could be too fast or too slow and those beings would probably never decode it. Therefore, it is of interest to device a way to encode messages agnostic of time scale. Fractal messaging would allow one to do this.
Fractals show up as interesting carriers for messages which need to be scale-independent, especially the ones with structural self-similarity or even scale invariance. By starting from a spatial embedding rationale, a framework will be developed for a time scale-free messaging alternative. Thus, in the present paper, a theoretical framework shall be developed to provide time-agnostic communication, hereby understood as the transmission of binary messages. A viable suggestion for message embedding and transmission based on software and hardware implementations will be discussed, as well as how the framework provides time scale-free messages.
2 Preliminaries
2.1 Fractals
According to Mandelbrot, a fractal is a set whose Hausdorff-Besicovitch dimension strictly exceeds its topological dimension [10]. A remarkable characteristic of fractals is that they can have a similar geometrical aspect at various scales, a behavior called self-similarity. Sometimes, if such aspect is exactly the same at every scale, the shape is called affine self-similar or even scale invariant [11, 12]. As discussed in the Introduction, fractals are interesting candidates to be carriers of signals, lending to those their self-similarities. The next subsections present two important fractals which are considered in the present paper.
2.1.1 The Mandelbrot set
The Mandelbrot set is a set of complex numbers (therefore two-dimensional if plotted in the complex plane) defined by a simple recursion formula that hides intricate patterns and great complexity. It can be defined as the set of complex numbers $c$ for which the function $f_{c}(z)=z^{2}+c$ remains finite after being infinitely iterated from $z=0$. This set was first defined and drawn by Brooks and Matelski [13] in 1978, but became famous in 1980 when Mandelbrot [14] obtained high-quality visualizations of the set.
The boundary of the Mandelbrot set, when plotted in the complex plane, produce infinitely complicated patters, especially when zoomed in. Such boundary is a fractal curve, following Mandelbrot‘s definition. Despite its simple definition, the fact that the set is bidimensional may pose challenges to the message embedding process. Thus, lower dimensional alternatives, such as the Weierstrass function, could be considered.
2.1.2 The Weierstrass function
Weierstrass defined a function which is continuous but nowhere differentiable as the following Fourier series:
$W_{a,b}(t)=\sum_{n=0}^{\infty}a^{n}\cos(b^{n}\pi t),$ (1)
where $10<a<$ 1, $b$ is a positive odd integer, $t\in\mathbb{R}$ and
$ab>1+\frac{3}{2}\pi.$ (2)
Hardy [15] further discussed these restrictions to show that whenever $0<a<1<b$ and $ab\geq 1$, for $b$ not necessarily integer, the function does not possess a finite differential coefficient at any point. This latter set of restrictions shall be considered in the present paper.
Figure 1 presents the general behavior of $W_{a,b}(t)$ as the number of terms (or components) in the summation increase.
Figure 1: General behavior of $W_{a,b}(t)$ as more terms (components) are considered in the summation when $a=1/2$ and $b=3$.
Figure 1 indicates that $W_{a,b}(t)$ is nothing but the superposition of sinusoidal waves, which have been extensively used as carriers for messages using traditional radio transmission techniques. In the next section, a specific technique for amplitude modulation is discussed.
2.2 Amplitude modulation of signals
There are several ways to encode binary messages by using carrier signals. In the present paper the so called Amplitude Shift Keying (ASK) scheme shall be considered, where one simply multiply the wave form of a certain carrier wave by another signal formed by the sum of several rectangle waves representing the bits.
Figure 2 presents a simple modulation of the signal $f(t)$. Considering that each bit modulates a single period of the signal $f(t)$, it can be seen the binary “10” was used to modulate it.
Figure 2: Simple illustration of amplitude modulation (or masking) of a signal $f(t)$.
Now that all the preliminary concepts have been presented, the general intuition behind fractal messaging can be discussed.
3 Intuition
Formally, one can define the recurrence relation used to build the Mandelbrot set as: Let $z_{0}=0$ and $z_{n+1}=z_{n}^{2}+c$, for a fixed complex number $c$. The points belonging to the set are the values of $c$ such that $\lim\limits_{n\to\infty}|z_{n}|<R_{e}$, where $R_{e}$ is the escape radius chosen for the algorithm. It is easy to prove using the triangle inequality that if $|z_{n}|>2$, then the sequence will diverge, implying that $R_{e}=2$ is a natural choice, although other values of $R_{e}\geq 2$ can be considered.
For the values not belonging to the Mandelbrot set, let $z_{e}$ represent the minimum value of $z_{n}$ such that $|z_{n-1}|<R_{e}$ and $|z_{n}|>R_{e}$, where $|w|$ is the absolute value of the complex number $w=w_{R}+iw_{I}$. In that case, $w_{R}=Re(w)$ and $w_{I}=Im(w)$ are the Real and Imaginary parts of $w$, respectively. Then, since the values of $z_{e}$ are unique for a given escape radius, it is possible to use them to perform color encodings of information into the complement of the Mandelbrot set in the complex plane. For example, for binary encodings, Figure 3 presents the plots obtained for $Re(z_{e})\geq 0$ and $Im(z_{e})\geq 0$, respectively. Both figures illustrate that one can embed messages, such as a binary counter representing all positive integers or natural numbers, in the complement of the Mandelbrot set, respecting the same properties of self-similarity and, therefore, invariance of the sender/receiver spatial scales.
Figure 3: Simple illustration of embedding $Re(z_{e})\geq 0$ (left) and $Im(z_{e})\geq 0$ (right) for $R_{e}=10$
One may notice that by considering the message to be transmitted as a bidimensional image, some sense of orientation is mandatory for the recipient to properly understand it, just like the case of the Arecibo message - which becomes an unintelligible set of characters if not arranged properly. This can be challenging to ensure in practice. Besides, the reproduction of images could rely on specific sensory characteristics (vision, for example), which are not granted to be among the receiver entity’s capacities. This has been explored by Vakoch [16] while addressing the challenge of communicating information unique to one civilization, highlighting that sending two-dimensional pictorial representations of mathematical concepts and physical objects may not be readily intelligible to a civilization with different conventions than ours.
On the other hand, Vakoch [16] indicated that the use of three-dimensional representations redundantly encoded in multiple formats (e.g., as both vectors and as rasters) could be better suited for the communication task. Overall, one of the core ideas presented is that, after a gradual process of presenting familiar objects (stars, for example), unfamiliar concepts could be introduced and that redundancy could increase the likelihood of intelligibility.
The general idea of redundancy as a means to attain intelligibility is deeply related to the use of fractals as carriers for messages, but from a transmission point of view rather than the content of the message itself. In this context, devising a way to transmit messages such that redundancy is achieved both spatially and temporally is of interest, especially if this could be implemented in a compact unidimensional spatial setting. This approach shall be described in the next section.
4 Using Weierstrass function’s components as carriers for binary messages
When a space and time-agnostic framework for message transmission is considered, it would be interesting to encode a message so that it could be decoded along several space and time scales. One shall consider that any “being” receiving the message would live in a specific time scale, and that messages could be “understood” as long as they were decoded and reproduced in suitable frequencies, for example.
This way, the core idea of the framework hereby proposed is to encode a binary message along infinitely many (power-like distributed) frequencies, transmit such message and then decode and reproduce it. Starting from uni-dimensional messages, which are simply streams of binary information, a natural carrier would be uni-dimensional electromagnetic waves. Thus, it is of interest to consider fractals which could be viewed as uni-dimensional electromagnetic waves. This is the case of the Weierstrass function and its components.
Figure 2 indicates how to modulate a single sinusoidal signal in order to embed a binary stream in it. For different types of receiver entities, each functioning on a given spatio-temporal scale, the message could only be perceived if reproduced as a suitable signal in terms of amplitude and frequency. Therefore, since the correct scales are unknown, a direct way to guess them is by repeating the message across the whole frequency-amplitude spectrum. In order to cover infinitely many frequency ranges, one can modulate each of the components of the summation in Eq. 1. In order to ensure a single-period single-bit modulation rationale, one must make the rectangular waves representing each bit match a single period of the carrier wave. This can be done by adjusting the frequencies and performing phase shifts on the rectangular waves. Mathematically, let $A_{m}(t,msg)$ be the $m$-th amplitude modulation function representing the same binary message $msg$, the resulting fractal communication signal $\mathcal{F}_{a,b}$ can be represented as:
$\mathcal{F}_{a,b}(msg,t)=\sum_{m=0}^{\infty}a^{m}A_{m}(t,msg)\cos(b^{m}\pi t),$ (3)
The series representing $\mathcal{F}_{a,b}$ is convergent whenever the restrictions for $a$ and $b$ presented for $W_{a,b}$ are satisfied, since:
$\mathcal{F}_{a,b}(msg,t)\leq W_{a,b}(t).$ (4)
By considering equation 3 as the signal to be transmitted, not only infinitely many frequencies are covered (controlled by $b$), but also infinitely many amplitudes (controlled by $a$).
4.1 Example of encoded messages in fractal carriers
As a simple example, one shall encode the message “hi!” using its ASCII binary representation ‘011010000110100100100001’. Considering positive time as our $t$ variables, Figure 4 presents the encoded fractal message.
Figure 4: Encoded message “hi!” using the components of $W_{1/2,3}$ as carriers. Ten components were considered, for computational simplicity.
Fractal behavior is observed for the encoded message by zooming in the lower plot in Figure 4. These zoomed in regions are presented in Figure 5.
Figure 5: Zooming in $\mathcal{F}_{1/2,3}(^{\prime}hi!^{\prime},t)$. Twenty components were considered, for computational simplicity.
It is interesting to notice that after embedding the message, the fractal dimension of the resulting signal changes. In the case of Figure 4, it was believed that $W_{a,b}$ had a fractal dimension of $2+\log_{b}(a)$ [11], being such result later proved for some values of $a$ and $b$ [17]. Thus, $W_{1/2,3}$ has a fractal dimension of $2-\log_{3}(2)\approx 1.36907$, which could be verified using a Box-counting algorithm. On the other hand, using the same algorithm, the fractal dimension of $\mathcal{F}_{1/2,3}(^{\prime}hi!^{\prime},t)$ was found to be approximately $1.42146$, slightly higher than the fractal dimension of $W_{1/2,3}$. This suggests that by embedding messages, even though the ASK scheme used actually suppresses information (by not transmitting the signal in the positions of the bit 00 in the message), the overall complexity of the signal increases as the fractal dimension can be considered its proxy [10].
4.2 Decoding and reproducing
It is possible to use inverse Fourier transform tools to retrieve the original message because the amplitude modulation behavior repeats itself periodically (infinite loop). In order to reproduce the message, either sounds (pressure waves) or laser pulses could be generated. One caveat of these approaches is fine-tuning the energy content of the reproduction. This is important to avoid under or over detection (the laser, for example, could be undetectable in a high energy - high frequency - scenario or could be as powerful as the solar system’s sun in a low energy - low frequency - scenario).
Another interesting approach is to use one-dimensional chains of trapped ions to produce a quantum mechanical system with discrete scale invariance and fractal-like time dependence [18]. In that case, it was shown that patterns similar to the Weierstrass function could be produced, which are pivotal to encode the messages using the framework hereby proposed. Considering the findings of Rose and Wright [19], sending such chains of trapped ions together with circuits to modulate the amplitude of the waves emitted could be an interesting transmission approach at hardware level. There are certainly limitations, but, in principle, for a certain range of amplitudes and frequencies, the transmission and emission could be achieved. Regarding software implementations, there are also limits up to which the waves can be generated and reproduced. These come from the limitations of precision for the calculations considered, for example.
4.3 Beyond binary encodings of characters: Iterated Function Systems (IFSs) and Turing machines as fractals
In the present paper, the messages being transmitted were binary streams representing, for example, ASCII encodings of known characters. By using fractal carriers, the message and the carriers became a single entity, which was shown to be a fractal. This suggests a more general framework where concepts could be stated in terms of fractals, which could ultimately encode not only characters, but also computation concepts in an integrated fractal way.
As a matter of fact, using fractals to establish a communication system may be beneficial as the universe has been shown to present fractal-like structures with remarkable self-similarity (the self-similar multifractal distributions of dark matter and gas, for example [20]). Therefore, the entities receiving the message could be familiar with self-similar structures, due to their occurrence in their own environment. Not only that, but literature indicates that digital sundials have been devised as projections of shadows of a fractal set [21], which would allow the possible communication of a correspondence between time scales related to Earth’s sun and the receivers’ sun. Therefore, using fractals as communication entities could be of interest.
Regarding possible relations between fractal geometry and computation, Penrose suggested that fractals might be a graphical way of looking at non-recursive mathematics [22]. Further exploring that rationale, Dube [23] established a relationship between the classical theory of computation and fractal geometry. In particular, he considered the concepts behind IFSs to define fractals, indicating that for every Turing machine (in the context of computation over integers and rationals) there exists a fractal set which can be viewed, in a certain sense, as geometrically encoding the complement of the language accepted by the machine. This suggests that in a fractal-based communication system, it would be possible to transmit fractal-based geometrical models of computation which are computationally universal. Inspired by previous works [21], Dube [23, 24] proposed a fractal-based “sun-computer”, indicating that for every Turing machine $M$ there exists such a fractal set $A$ which graphically “encodes” the behavior of $M$ on all inputs. This could be a potential carrier for a Turing Machine computational model without the need to use binary encodings.
5 Conclusion
We have introduced a new mathematical framework based on fractal theory that allows scale-independent communication.
The approach instantiates the idea of sending a self-encoding and self-decoding signal as a mathematical formula that carries both the data and its unfolding instructions. This would be similar to sending a self-executable message with its computer to run it. It is equivalent to sending self-executable computer code that, when unfolded, reads a message in all possible time scales and at the same time as if in all possible channels simultaneously.
Together with our previous work [5], we have contributed to specific technical challenges at the receiver and sender ends.